• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name pking

    Log entry time 08:02:42 on October 19, 2009

    Entry number 295649

    keyword=Owl shift summary

    SL : Paul King
    TO: Richard Wilson
    3rd: Joshua G. Rubin


    Production runs at 98 uA with IHWP IN:
    	13388	45.7
    	13389	43.9
    	13390	43.7
    	13391	44.3
    	13392	40.5
    	13393	43.0
    	13394	35.8
    	13395		Mostly beam off
    	13396	41.6


    Shift starts with HAPPEX run 13388 and Compton run 20500 in progress.

    00:17---Compton HV trips. HV trips again at 00:22. Again at 00:27.

    00:32---Start new HAPPEX run, 13389.

    00:33---Compton ROC6 stops responding. Power cycle ROC6 with hareboot4 "outlet 6". Do "coda reboot"

    Compton HV trips at 00:45. Compton HV trips again at 01:02. Again at 01:07. Again at 01:13.

    00:14---Compton runcontrol screen just disppeared. Session closed from adaql3 to the compton computer. Restarted runcontrol and started a new Compton run.

    Compton HV trips at 01:17. Again at 01:21. Again at 01:26.

    01:29---Start new HAPPEX run, 13390.

    01:32---D.Wang takes over the Compton DAQ.

    01:42---Compton HV trips again. There was not a run in progress, so we can't see the rates in the archiver.

    01:52---Take run 25079 to check spot++.

    02:25---Start new HAPPEX run, 13391.

    03:22---Start new HAPPEX run, 13392.

    03:29---D.Wang is finished with his tests on the Compton DAQ. Begin Compton run 20508.

    03:46---Compton HV tripped.

    04:15---Compton HV tripped. Tripped again at 04:21. Tripped again at 04:30. Tripped again at 04:32. Tripped again at 04:40.

    04:41---Start new HAPPEX run, 13393.

    04:42---Take run 25080 for spot++ check.

    04:49---Compton HV tripped again at 04:49. Tripped again at 04:58. Tripped again at 05:08. Tripped again at 05:12. Tripped again at 05:13.

    05:16---Start new HAPPEX run, 13394.

    05:16---Compton HV tripped again.

    05:18---Started new Compton run, 20509.

    05:22---Compton HV tripped again at 05:22. Tripped again at 05:24. Tripped again at 05:25.

    OPS hasn't found any smoking gun for why we're getting the rate spikes. Sometimes they are associated with energy drifts, but not always.

    Tripped again at 05:31. Tripped again at 05:32. Tripped again at 05:34. Tripped again at 05:37 Tripped again at 05:41. Tripped again at 05:43. Tripped again at 05:46. Tripped again at 05:47. Tripped again at 05:48. Tripped again at 05:49. Tripped again at 05:51. Tripped again at 05:53. Tripped again at 05:55. Tripped again at 05:58.

    06:00---Call out to Alex.

    Changed Compton HV to -1750. Move the table to X=53.50 Reduce HV to -1600 V. Reduce HV to -1450 V. Change attenuator to 1 and gain to 8.

    OPS will try a pathlength check.

    Alex will come in soon. If it trips again, leave it off until he gets in.

    06:13---Start new HAPPEX run, 13395. Mostly no beam in this run.

    06:41---OPS is done with the pathlength adjustment Sending beam back to us now.

    Start new HAPPEX run, 13396.

    06:50---The Compton rates are not stable, and don't show any difference between laser on and off. Alex was expecting to come in shortly.

    06:57---Noticed the beam current was right about 96 uA. Request it be increased a little to be about 98 uA.

    07:15---Alex arrives and starts to look at the Compton.

    07:39---Start new HAPPEX run, 13397.